OpenStudy (keke):

simplify; √(5) over 7-√(10)

OpenStudy (anonymous):

\[\frac{\sqrt{5}}{7-\sqrt{10}}\times \frac{7+\sqrt{10}}{7+\sqrt{10}}\]

OpenStudy (anonymous):

\[7\sqrt{5} + 5\sqrt{2}\]

OpenStudy (anonymous):

upon 39

OpenStudy (anonymous):

\[7\sqrt{5} + 5\sqrt{2}\] ---------------------- 39